Problem 1 - Unit Conversions A race car moves on a straight line with a constant velocity v 300 km/h; rewrite v in m/s and mph. (10 points 998 kg/m3; what is the density of The density of water in a glass in this room is about ρ water in g/em3? (10 points)
Answer #1

Solution:

a) v = 300 km/h

In meter per second

=> v = 300 * 1000 m / 60 * 60 sec

=> v = 83.33 m/s

In miles per hour

and , v = 300 km/h

=> v = 300 * 0.62137 miles / h

=> v = 186.4 miles/h

b) p = 998 kg / m^3

=> p = 998 * 1000 / 10^6

=> p = 998 * 10^-3 g / cm^3

=> p = 0.998 g / cm^3

please rate thanks.